第48回環論および表現論シンポジウム（2015年）第二報

第 48 回環論および表現論シンポジウムのプログラムが出来ましたのでお知らせいたします。また懇親会もあわせてご案内申し上げます。このシンポジウムは以下の研究費の援助のもとに行われます。

日本学術振興会科学研究費補助金基盤研究 (B) 研究代表者：浅芝秀人（静岡大学）研究代表者：伊山修（名古屋大学）日本学術振興会科学研究費補助金基盤研究 (C) 研究代表者：馬場良始（大阪教育大学）研究代表者：池畑秀一（岡山大学）研究代表者：西中恒和（兵庫県立大学）研究代表者：荒谷督司（岡山理科大学）日本学術振興会海外共同研究・セミナー（日中パートナーシップ）研究代表者：佐藤眞久（山梨大学）

プログラム責任者西中恒和（兵庫県立大学）

記

日程 : 2015 年 9 月 7 日 (月)〜10 日 (木)

場所 : 名古屋大学 多元数理科学棟 509 室

〒 464-8601 愛知県名古屋市千種区不老町

       http://www.math.nagoya-u.ac.jp/ja/direction/nagoya-st.html
       http://www.math.nagoya-u.ac.jp/ja/direction/campus.html

会場責任者 : 伊山修 (名古屋大学 iyama@math.nagoya-u.ac.jp)

プログラム責任者 :西中恒和（兵庫県立大学 nishinaka@econ.u-hyogo.ac.jp )

懇親会 : 日時 : 9 月 9 日 (水) 18:30 〜会場 : 名大生協「レストラン花の木」（名古屋大学キャンパス内）会費 : 5,000 円 (学生 4,000 円）

懇親会参加希望者は 8 月 21 日 (金) までに、 E-mail にて、会場責任者・伊山修氏 (iyama@math.nagoya-u.ac.jp) までお知らせ下さい。多数の方のご参加をお待ちしています。

9 月 7 日（月）

9:00–9:30 小原大樹（東京理科大学） On the Hochschild cohomology ring modulo nilpotence of the quiver algebra with quantum-like relations

9:40–10:10 亀山統胤（信州大学）・星野光男（筑波大学）・古賀寛尚（東京電機大学）

Crossed products for matrix rings 10:20–10:50 浅芝秀人（静岡大学）

Cohen-Montgomery duality of bimodules and its applications to equivalences given by bimodules

11:00–11:50 Jan Schröer (University of Bonn)

Convolution algebras and enveloping algebras I 13:40–14:10 劉裕（名古屋大学）

Cotorsion pairs on triangulated and exact categories 14:20–14:50 水野有哉（名古屋大学）・山浦浩太（山梨大学）

Higher APR tilting preserve n-representation infiniteness 15:00–15:30 毛利出（静岡大学）

2-dimensional quantum Beilinson algebras 15:50–16:40 有木進（大阪大学）

Survey on representation theory of quiver Hecke algebras I 16:50–17:40 Steffen Koenig （Universität Stuttgart）

Some special cases of Morita-Tachikawa correspondence

9 月 8 日（火）

9:00–9:30 小境雄太（東京理科大学）・功刀直子（東京理科大学）

Construction of two-sided tilting complexes for Brauer tree algebras 9:40–10:10 足立崇英（名古屋大学）

The classification of two-term tilting complexes for Brauer graph algebras 10:20–10:50 古賀寛尚（東京電機大学）・星野光男（筑波大学）

Derived equivalences and Gorenstein projective dimension 11:00–11:50 Jan Schröer (University of Bonn)

Convolution algebras and enveloping algebras II 13:40–14:10 吉田健一（日本大学）・後藤四郎（明治大学）・谷口直樹（明治大学）・松岡直之（明治大学）

Almost Gorenstein Rees algebras 14:20–14:50 松井紘樹（名古屋大学）

The structure of preenvelopes with respect to maximal Cohen-Macaulay modules

15:00–15:30 高橋亮（名古屋大学）

Thick subcategories of derived categories of isolated singularities 15:50–16:40 有木進（大阪大学）

Survey on representation theory of quiver Hecke algebras II 16:50–17:40 綿谷安男（九州大学）

Quivers, operators on Hilbert spaces and operator algebras I

9 月 9 日（水）

9:00–9:30 山中聡（岡山大学）

On weakly separable extensions and weakly quasi-separable extensions 9:40–10:10 植田玲（島根大学）・丸林英俊（徳島文理大学）

Examples of Ore extensions which are maximal orders whose based rings are not maximal orders

10:20–10:50 Yotsanan Meemark （Chulalongkorn University）

Cayley Graphs over a Finite Chain Ring and GCD-graphs 11:00–11:50 綿谷安男（九州大学）

Quivers, operators on Hilbert spaces and operator algebras II 13:40–14:10 神田遼（名古屋大学）

Atom-molecule correspondence in Grothendieck categories 14:20–14:50 板場綾子（静岡大学）

Finite condition (Fg) for self-injective Koszul algebras 15:00–15:30 源泰幸（大阪府立大学）

Higher products on Yoneda Ext algebras 15:50–16:40 Steffen Koenig （Universität Stuttgart）

Filtered categories and representations of boxes 16:50–17:40 佐藤眞久（山梨大学）

特別講演： Report on the 7th China-Japan-Korea International Conference on Ring Theory

18:30– 懇親会

9 月 10 日（木）

9:00–9:30 淺井聡太（名古屋大学）

The Grothendieck groups of mesh algebras 9:40–10:10 平松直哉（呉工業高専）

Stable degenerations of Cohen-Macaulay modules over simple singularities of type (A_{n})

10:20–10:50 木村雄太（名古屋大学）

Tilting theory of preprojective algebras and c-sortable elements 11:10–11:40 上山健太（弘前大学）毛利出（静岡大学）

Tilting objects for noncommutative quotient singularities